OpenStudy (anonymous):

Use algebra to evaluate the limit. limh→ 0[(2(5 + h)^2 − 50)/h]

OpenStudy (tkhunny):

Well, do it! Expand (5+h)^2 and you'll nearly be done.

OpenStudy (debbieg):

Expand the binomial in the numerator, simplify, and then you'll be able to just substitute in the h=0.

OpenStudy (anonymous):

I got -10 and it said it was incorrect.

OpenStudy (debbieg):

Yeah, that's not it. Can you show your work? what did you get when you expanded the numerator?

OpenStudy (debbieg):

(2(5 + h)^2 − 50) = ?

OpenStudy (anonymous):

i got 20h+2h^2

OpenStudy (debbieg):

good. Now divide that by h. What do you get?

OpenStudy (anonymous):

20+2h

OpenStudy (debbieg):

Right! So you have: [(2(5 + h)^2 − 50)/h] =20+2h Now let h->0. What do you get?

OpenStudy (anonymous):

Oh, i see what i did! Thank you!

OpenStudy (debbieg):

You're welcome! :)